OpenStudy (anonymous):

You have a part time job. you work for 3h on Friday and 6h on Saturday. You also receive an allowance of $20 per week. You earn $92 per week. How much do you earn per hour at your part time job?

OpenStudy (anonymous):

1. you add 3 and 6 and divide it by 92 2. 9/92 = 1.2 = 1.20 3. ps dont add the 20$

OpenStudy (anonymous):

(3x+6x) X=how much you get paid an hour. 9x+20=92 subtract 20 from both side and solve for X